OpenStudy (anonymous):

f(x) = x2 - 81. Find f-1(x)

OpenStudy (whpalmer4):

\[y=f(x) = x^2-81\] To find \(f^{-1}(x)\), swap the variables and solve for \(y\): \[x = y^2-81\]solve for \(y\)

OpenStudy (anonymous):

We are having function as \[f(x)=x^2-81\] To find inverse of given function first replace f(x) by y

OpenStudy (whpalmer4):

Note that \(f^{-1}(x)\) is not a function, because it has multiple values for some values of \(x\).

OpenStudy (whpalmer4):

Also note that \(f(x)\) and \(f^{-1}(x)\) are always copies of each other reflected across the line \(y = x\)

OpenStudy (anonymous):

\[f(x)=x^2-81\] After replacing f(x) by y then we get \[y =x^2-81\] Now interchange roles of x and y

OpenStudy (anonymous):

Hello are you getting this ??

OpenStudy (anonymous):

After interchanging roles of x and y we get \[x=y^2-81\] Now try to solve this for y